OpenStudy (anonymous):

How much Zakat should be paid on a yearly savings of Rs 93,800?

OpenStudy (paki):

In Islam zakat is 2(1/2)% per year... so it makes [ 2(1/2) / 100 ] = [ 5/2 /100 ] = 5/200... so take (5/200)*93800=X... solve for "x" now... @sheelo_mughal

OpenStudy (anonymous):

can u explain me in detail?

OpenStudy (paki):

bro i have already explained it in detail..... what you want to ask... tell me now... sorry for late reply... actually OS was not working....

OpenStudy (paki):

@sheelo_mughal